OpenStudy (anonymous):

what is the simplified form of 15x^8/24y^5 divided by 4x^4/8y^2

OpenStudy (anonymous):

15x^8/24y^5 divided by 4x^4/8y^2 = (15x^8/24y^5 )(8y^2 /4x^4) consider x first x^8 /x^4 =x x x x x x x x / x x x x =x x x x =x^(8-4) =x^4 similarly, for y y^2 / y^5 =1/y^3 for number =15/24 8/4 =5/4 so ans =(5/4)(x^4)(1/y^3) =5x^4 / 4y^3

OpenStudy (anonymous):

thanks :D